grams of 1.0ft3 of Hg


1
Expert's answer
2021-05-10T01:51:29-0400

Density of mercury "=" "13.6g{cm}^{-3}"


Density of mercury in "g{ft}^{-3}=\dfrac{13.6}{(0.0328)^3}" "=3.86\times10^5g{ft}^{-3}"


Now


Density("g{ft}^{-3}" )"=\dfrac{mass(gm)}{volume({ft}^{3})}"


"3.86\times10^5=\dfrac{mass(gm)}{1}"


So mass in grams of "1.0ft^3" of Hg is "3.86\times10^5" gm